OpenStudy (anonymous):

A parallelogram has the vertices (-1, 2), (4, 4), (2, -1) and (-3, -3). Determine what type of parallelogram [10 points]. Find the perimeter and area [20 points].

OpenStudy (anonymous):

@DSS could you help again?

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

give me a min

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what do you mean by " type of parallelogram " ??

OpenStudy (anonymous):

what is the shape

OpenStudy (anonymous):

@Akhilender

OpenStudy (anonymous):

Using your distance formula, you see that all sides of the parallelogram are the same length, which is \[\sqrt{29}\]

OpenStudy (anonymous):

Perimeter is s+s+s+s. so \[\sqrt{29}+\sqrt{29}+\sqrt{29}+\sqrt{29}=4\sqrt{29}\] units

OpenStudy (anonymous):

A=s*s\[\sqrt{29}*\sqrt{29}=29 \square units\]

OpenStudy (anonymous):

I don't know if this is correct but you've got a parallelogram with 4 equal sides...I'd say a rhombus. The angles aren't equal so you can't have a square.

OpenStudy (anonymous):

thank you so much! you honestly have no idea how much this helps me

OpenStudy (anonymous):